There were​ 36,000 people at a ballgame in Los Angeles. The​ day's receipts from​ $12 reserved seats and​ $6 general-admission seats were​ $258,000 How many of each type of seat were​ sold?
Answer by ikleyn(52780)

.
There were​ 36,000 people at a ballgame in Los Angeles. The​ day's receipts from​ $12 reserved seats and​
$6 general-admission seats were​ $258,000 How many of each type of seat were​ sold?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Let x = # of $12; y = # of $6.
Then

  x +  y =  36000,   (1)
12x + 6y = 258000.   (2)

Or, after canceling the factor 6 in (2)

 x + y = 36000,      (1')
2x + y = 43000.      (2')

Now distract equation (1') from (2').